Calculate the depth in the ocean at which the
pressure is three times atmospheric pressure.
Atmospheric pressure is 1.013 × 105 Pa. The
acceleration of gravity is 9.81 m/s
2
and the
density of sea water is 1025 kg/m3
.
Answer in units of m.
1
Expert's answer
2020-02-28T10:01:44-0500

"p=p_0+\rho gh"




"3p_0=p_0+\rho gh \to h=\frac {2p_0}{\rho h}=\frac{2\cdot101325}{1025\cdot 9.81}=20.1m"